# Kennedy-Tasaki transformation and non-invertible symmetry in lattice models beyond one dimension

### v1

@@ B.2 Two dimension
$\textbf{D}_{r}^{(2)^{\prime}}$ $\equiv\mathbf{\tilde{P}}^{\prime}\tilde{\mathbf{D}}_{x}H\tilde{\mathbf{D}}_{y}\mathbf{\tilde{P}},$ (166)
@@
$\begin{split}\mathbf{\tilde{P}}^{\prime}&\equiv\prod_{l=1}^{j}\frac{\left(1-\eta^{x}_{l}\right)}{2}\prod_{l=j+1}^{L_{y}}\frac{\left(1+\eta^{x}_{l}\right)}{2}\prod_{k\neq 1,i,i+1}\frac{\left(1+\eta^{y}_{k}\right)}{2}\frac{(1-\eta^{y}_{i})}{2}\frac{(1-\eta^{y}_{i+1})}{2}\frac{(1+(-1)^{j}\eta^{y}_{1})}{2},\end{split}$ (167a)
$\mathbf{\tilde{P}}$ $\equiv\prod_{l\neq j}\frac{(1+\eta^{x}_{l})}{2}\prod_{k\neq i}\frac{(1+\eta^{y}_{k})}{2}\frac{(1-\eta^{x}_{j})}{2}\frac{(1-\eta^{y}_{i})}{2},$ (167b)
@@
Furthermore our mapping of the Pauli $X$ operator is shifted diagonally to the other sublattice. We also find the composition of two $\overline{\text{KT}}^{(2)}$ transformations gives
$(\overline{\text{KT}}^{(2)})^{2}=\mathbf{T}_{r}\mathbf{T}_{b}\mathbf{P}_{r}\mathbf{P}_{b},$ (168)
where $\mathbf{T}_{r}$ and $\mathbf{T}_{b}$ are diagonal translations on the red and blue sublattices, respectively. $\mathbf{P}_{r}$ and $\mathbf{P}_{b}$ are the respective projections onto the Hilbert spaces of the red and blue sublattices, which are separately subsystem symmetric.
@@ B.3 Higher dimension
where the term in the L.H.S. and R.H.S. are related by diagonal translation by half unit. By $r\leftrightarrow b$ we get another similar set of equations. Composition of two $\overline{\text{KT}}^{(d)}$ transformations is
$(\overline{\text{KT}}^{(d)})^{2}=\mathbf{T}_{r}^{(d)}\mathbf{T}_{b}^{(d)}\mathbf{P}_{r}^{(d)}\mathbf{P}_{b}^{(d)},$ (170)
where $\mathbf{T}_{r}^{(d)}$ and $\mathbf{T}_{b}^{(d)}$ are diagonal translations on the red and blue sublattices. $\mathbf{P}_{r}^{(d)}$ and $\mathbf{P}_{b}^{(d)}$ are projections onto subsystem symmetric subspace of the Hilbert spaces of red and blue sublattices.
@@ C.1 $\mathbb{Z}_{N}$ plaquette transverse field clock model
$H_{TFPC}=-\sum_{i,j}\left(Z_{i,j}Z^{\dagger}_{i+1,j}Z^{\dagger}_{i,j+1}Z_{i+1,j+1}+\text{h.c}+\lambda(X_{i,j}+X_{i,j}^{\dagger})\right).$ (171)
@@
$\tilde{\mathbf{D}}_{(N),x}\equiv\prod_{j=1}^{L_{y}}D_{(N),j}^{x},\qquad\tilde{\mathbf{D}}_{(N),y}\equiv\prod_{j=1}^{L_{x}}D_{(N),j}^{y},$ (173)
where $\mathbf{D}_{(N),j}^{x}$ represent the operator $\mathbf{D}_{(N)}$ for a fixed horizontal row. Similarly $\mathbf{D}_{(N),j}^{y}$ represent the operator $\mathbf{D}_{(N)}$ for a fixed vertical column. We define the following operator, a quantum Fourier transform that is a generalization of the Hadamard gate to qudit d.o.f.,
$H_{(N)}\equiv\frac{1}{\sqrt{N}}\sum_{x=0}^{N-1}\sum_{y=0}^{N-1}e^{\frac{2\pi ixy}{N}}\ket{x}\bra{y}.$ | . (174)
@@
$H_{(N)}Z=X^{\dagger}H_{(N)},\qquad H_{(N)}X=ZH_{(N)}.$ (175)
@@
Note that the operator $H_{(N)}$ for $N=2$ is the same as the Hadamard transformation. We define the operator
@@
$\tilde{\mathbf{D}}_{(N)}\equiv\tilde{\mathbf{D}}_{(N),x}H_{(N)}\tilde{\mathbf{D}}_{(N),y},\quad\mathbf{P}_{(N)}\equiv\prod_{j=1}^{L_{y}}\frac{(1+\eta^{x}_{j}+(\eta^{x}_{j})^{2}+...+(\eta^{x}_{j})^{N-1})}{N}\prod_{i=1}^{L_{x}}\frac{(1+\eta^{y}_{i}+(\eta^{y}_{i})^{2}+...+(\eta^{y}_{i})^{N-1})}{N}.$ (176)
The non-invertible KW duality operator we define is $\mathbf{D}_{(N)}=\mathbf{P}_{(N)}\tilde{\mathbf{D}}_{(N)}\mathbf{P}_{(N)}$. The following relations are satisfied by $\mathbf{D}_{(N)}$,
$\mathbf{D}_{(N)}X_{i,j}=Z_{i,j}Z^{\dagger}_{i+1,j}Z^{\dagger}_{i,j+1}Z_{i+1,j+1}\mathbf{D}_{(N)},$ (177a)
$\mathbf{D}_{(N)}Z_{i,j}Z^{\dagger}_{i+1,j}Z^{\dagger}_{i,j+1}Z_{i+1,j+1}=X^{\dagger}_{i+1,j+1}\mathbf{D}_{(N)}.$ (177b)
@@
The operator $\mathbf{D}_{(N)}$ commutes with the Hamiltonian at $\lambda=1$. After an appropriate phase normalization for $\mathbf{D}_{(N)}$, we also have $\mathbf{D}_{(N)}^{2}=\mathbf{C}\mathbf{T}_{diag}\mathbf{P}_{(N)}$ where $\mathbf{T}_{diag}$ is the diagonal translation operator which send $(i,j)$ to $(i+1,j+1)$ and $\mathbf{C}$ is the conjugation operator with action $CX=X^{\dagger}C$ and $CZ=Z^{\dagger}C$.
C.2 Higher dimensional hypercubic clock models
@@
We generalize the discussion for the hypercubic clock model with transverse field to all spatial dimension. The Hamiltonian for the transverse field hypercubic clock model is given by
@@
$\mathbf{D}_{(N)}^{(d)}\equiv\mathbf{P}_{(N)}\tilde{\mathbf{D}}_{(N),d}\mathbf{H}_{(N)}\tilde{\mathbf{D}}_{(N),d-1}\mathbf{H}_{(N)}...\tilde{\mathbf{D}}_{(N),1}\mathbf{P}_{(N)},$ (179)
where each of the $\tilde{\mathbf{D}}_{(N),i}$ are string operators like Eq. (2) on straight line along $i^{\text{th}}$ direction and $\mathbf{P}_{(N)}$ is a projection onto subsystem straight line like symmetries. The operator $\mathbf{D}_{(N)}^{(d)}$ commutes with the Hamiltonian at $\lambda=1$. After an appropriate phase normalization for $\mathbf{D}_{(N)}^{(d)}$, acting the operator twice will generate a diagonal unit shift and a conjugation in d dimensions, i.e, $(\mathbf{D}_{(N)}^{(d)})^{2}=\mathbf{C}\mathbf{T}_{(N),(1,...,1)}\mathbf{P}_{(N)}$ where $\mathbf{T}_{(N),(1,...,1)}$ is a diagonal translation on the hypercube and $\mathbf{C}$ is a conjugation.

### v2

@@ B.2 Two dimensions
$\textbf{D}_{r}^{(2)^{\prime}}$ $\equiv\mathbf{\tilde{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}^{\prime}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}}\tilde{\mathbf{D}}_{x}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(2)}_{r}}\tilde{\mathbf{D}}_{y}\mathbf{\tilde{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}},$ (159)
@@
$\begin{split}\mathbf{\tilde{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}^{\prime}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}}&\equiv\prod_{l=1}^{j}\frac{\left(1-\eta^{x}_{l}\right)}{2}\prod_{l=j+1}^{L_{y}}\frac{\left(1+\eta^{x}_{l}\right)}{2}\prod_{k\neq 1,i,i+1}\frac{\left(1+\eta^{y}_{k}\right)}{2}\frac{(1-\eta^{y}_{i})}{2}\frac{(1-\eta^{y}_{i+1})}{2}\frac{(1+(-1)^{j}\eta^{y}_{1})}{2},\end{split}$ (160a)
$\mathbf{\tilde{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}}$ $\equiv\prod_{l\neq j}\frac{(1+\eta^{x}_{l})}{2}\prod_{k\neq i}\frac{(1+\eta^{y}_{k})}{2}\frac{(1-\eta^{x}_{j})}{2}\frac{(1-\eta^{y}_{i})}{2},$ (160b)
@@
Furthermore, our mapping of the Pauli $X$ operator is shifted diagonally to the other sublattice. We also find the composition of two $\overline{\text{KT}}^{(2)}$ transformations gives
$(\overline{\text{KT}}^{(2)})^{2}\propto\,\mathbf{T}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1,1),}r}\mathbf{T}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1,1),}b}\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{r}\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{b},$ (161)
where $\mathbf{T}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1,1),}r}$ and $\mathbf{T}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1,1),}b}$ are diagonal translations on the red and blue sublattices, respectively. $\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{r}$ and $\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{b}$ are the respective projections onto the Hilbert spaces of the red and blue sublattices, which are separately subsystem symmetric.
@@ B.3 Higher dimensions
where the terms in the L.H.S. and R.H.S. are related by diagonal translation by half unit. By $r\leftrightarrow b$ we get another similar set of equations. Composition of two $\overline{\text{KT}}^{(d)}$ transformations is
$(\overline{\text{KT}}^{(d)})^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\propto}\,\mathbf{T}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1,...,1),}r}\mathbf{T}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1,...,1),}b}\mathbf{P}_{r}^{(d)}\mathbf{P}_{b}^{(d)},$ (163)
where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{T}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1,...,1),}r}}$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{T}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1,...,1),}b}}$ are diagonal translations on the red and blue sublattices. $\mathbf{P}_{r}^{(d)}$ and $\mathbf{P}_{b}^{(d)}$ are projections onto subsystem symmetric subspace of the Hilbert spaces of red and blue sublattices.
@@ C.1 $\mathbb{Z}_{N}$ plaquette transverse field clock model
$\begin{split}H_{TFPC}&=-\sum_{i,j}\left(Z_{i,j}Z^{\dagger}_{i+1,j}Z^{\dagger}_{i,j+1}Z_{i+1,j+1}+\text{h.c}\right.\\ &\left.\hskip 85.35826pt+\lambda(X_{i,j}+X_{i,j}^{\dagger})\right).\end{split}$ (164)
@@
$\tilde{\mathbf{D}}_{(N),x}\equiv\prod_{j=1}^{L_{y}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{\mathbf{D}}_{(N),j}^{x}},\qquad\tilde{\mathbf{D}}_{(N),y}\equiv\prod_{j=1}^{L_{x}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{\mathbf{D}}_{(N),j}^{y}},$ (166)
where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{\mathbf{D}}_{(N),j}^{x}}$ represent the operator ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{\mathbf{D}}_{(N)}}$ for a fixed horizontal row. Similarly ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{\mathbf{D}}_{(N),j}^{y}}$ represent the operator ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{\mathbf{D}}_{(N)}}$ for a fixed vertical column. We define the following operator, a quantum Fourier transform that is a generalization of the Hadamard gate to qudit d.o.f.,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}}_{(N)}\equiv\frac{1}{\sqrt{N}}\sum_{x=0}^{N-1}\sum_{y=0}^{N-1}e^{\frac{2\pi ixy}{N}}\ket{x}\bra{y}.$ | . (167)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}}_{(N)}Z=X^{\dagger}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}}_{(N)},\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}}_{(N)}X=Z{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}}_{(N)}.$ (168)
@@
Note that the operator $\mathbf{H}_{(N)}$ for $N=2$ is the same as the Hadamard transformation. We define the operator
@@
$\tilde{\mathbf{D}}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}\equiv\tilde{\mathbf{D}}_{(N),x}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(2)}_{(N)}}\tilde{\mathbf{D}}_{(N),y},\quad\mathbf{P}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}\equiv\prod_{j=1}^{L_{y}}\frac{(1+\eta^{x}_{j}+(\eta^{x}_{j})^{2}+...+(\eta^{x}_{j})^{N-1})}{N}\prod_{i=1}^{L_{x}}\frac{(1+\eta^{y}_{i}+(\eta^{y}_{i})^{2}+...+(\eta^{y}_{i})^{N-1})}{N}.$ (169)
where $\mathbf{H}^{\otimes(2)}_{(N)}$ represent the tensor product of $\mathbf{H}_{(N)}$ on all sites of the two-dimensional lattice. The non-invertible KW duality operator we define is $\mathbf{D}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}_{(N)}=\mathbf{P}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}\tilde{\mathbf{D}}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}\mathbf{P}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}$. The following relations are satisfied by $\mathbf{D}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}$,
$\mathbf{D}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}X_{i,j}=Z_{i,j}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}Z_{i+1,j}Z_{i,j+1}Z_{i+1,j+1}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}\mathbf{D}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}},$ (170a)
$\mathbf{D}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}Z_{i,j}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}Z_{i+1,j}Z_{i,j+1}Z_{i+1,j+1}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}=X^{\dagger}_{i+1,j+1}\mathbf{D}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}.$ (170b)
@@
The operator $\mathbf{D}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}$ commutes with the Hamiltonian at $\lambda=1$. We also have $(\mathbf{D}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}})^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\propto}\,\mathbf{C}\mathbf{T}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1,1)}}\mathbf{P}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}$ where $\mathbf{T}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(1,1)}}$ is the diagonal translation operator which send $(i,j)$ to $(i+1,j+1)$ and $\mathbf{C}$ is the conjugation operator with action ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{C}}X=X^{\dagger}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{C}}$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{C}}Z=Z^{\dagger}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{C}}$.
C.2 higher-dimensional hypercubic clock models
@@
We generalize the discussion for the hypercubic clock model with a transverse field to all spatial dimensions. The Hamiltonian for the transverse field hypercubic clock model is given by
@@
$\mathbf{D}_{(N)}^{(d)}\equiv\mathbf{P}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(d)}}\tilde{\mathbf{D}}_{(N),d}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(d)}_{(N)}}\tilde{\mathbf{D}}_{(N),d-1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(d)}_{(N)}}...\tilde{\mathbf{D}}_{(N),1}\mathbf{P}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(d)}},$ (172)
where each of the $\tilde{\mathbf{D}}_{(N),i}$ are string operators like Eq. (2) on straight line along $i^{\text{th}}$ direction , ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(d)}_{(N)}}$ represent the tensor product of $\mathbf{H}_{(N)}$ on all sites of the $d$ dimensional lattice and $\mathbf{P}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(d)}}$ is a projection onto subsystem straight line like symmetries. The operator $\mathbf{D}_{(N)}^{(d)}$ commutes with the Hamiltonian at $\lambda=1$. Up to an overall normalization for $\mathbf{D}_{(N)}^{(d)}$, acting the operator twice will generate a diagonal unit shift and a conjugation in d dimensions, i.e, $(\mathbf{D}_{(N)}^{(d)})^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\propto}\,\mathbf{C}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{T}_{(1,...,1)}}\mathbf{P}_{(N)}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(d)}}$ where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{T}_{(1,...,1)}}$ is a diagonal translation on the hypercube and $\mathbf{C}$ is a conjugation.